Problem 4 The income-expenditure model draws a distinction between actual aggregate expenditure (AE) which is what is measured in the national accounts- and planned aggregate expenditure PAE). Use the following information to calculate AE and PAE Exports (X)20 Imports (M) 15 Consumption (C) 50 Planned Investment (IP) 25 Planned change in inventories-5 Unplanned change in inventories10 -Government spending (G) = 10 AE PAE
Answer #1

Equilibrium:

PAE = Total output or real GDP

The following equation is required for PAE.

PAE = Consumption + Planned investment + Government spending + (Export – Import)

         = 50 + 25 + 10 + 5; [as (X – M) = (20 – 15) = 5]

         = 90 (Answer)

AE is the addition of unplanned change in inventories and PAE. The following equation is required for AE

AE = PAE + Unplanned change in inventories

      = 90 + (- 10)

      = 90 – 10

      = 80 (Answer)

Note: planned change in inventories should not be considered here, since it is already considered in planned investment.